OpenStudy (anonymous):

The size of a certain cell is 2.5 x 10^-9m. Another cell is 1.5 x 10^3 times larger. How large is the larger cell in scientific notation? (1point) O 4 x 10^-27m O 4 x 10^-12m O 3.75 x 10^6m O 3.75 x 10^-6m Please Help!!!

OpenStudy (anonymous):

its C i believe

OpenStudy (anonymous):

Can you show work plz? @AsiaJenell

OpenStudy (anonymous):

Sure, hold on

OpenStudy (anonymous):

(2.5 * 10 ^-9)(1.5 * 10^3)= 2.5*1.5 X 10^(-9+3)= 3.75 X 10^6M